### v1

# MCITlib: Multimodal Continual Instruction Tuning Library and Benchmark

Abstract
Continual learning aims to equip AI systems with the ability to continuously acquire and adapt to new knowledge without forgetting previously learned information, similar to human learning. While traditional continual learning methods focusing on unimodal tasks have achieved notable success, the emergence of Multimodal Large Language Models has brought increasing attention to Multimodal Continual Learning tasks involving multiple modalities, such as vision and language. In this setting, models are expected to not only mitigate catastrophic forgetting but also handle the challenges posed by cross-modal interactions and coordination. To facilitate research in this direction, we introduce MCITlib, a comprehensive and constantly evolving code library for continual instruction tuning of Multimodal Large Language Models. In MCITlib, we have currently implemented 8 representative algorithms for Multimodal Continual Instruction Tuning and systematically evaluated them on 2 carefully selected benchmarks. MCITlib will be continuously updated to reflect advances in the Multimodal Continual Learning field. The codebase is released at https://github.com/Ghy0501/MCITlib.
1 Introduction
Enabling models to continuously learn and evolve like humans remains a fundamental challenge that limits the practical deployment of artificial intelligence systems in real-world scenarios. This difficulty primarily arises because models inevitably forget previously acquired knowledge when learning new information—a phenomenon known as catastrophic forgetting [32, 5, 21]. To mitigate this problem, a family of approaches collectively referred to as continual learning has been proposed. Traditional continual learning methods typically focus on unimodal tasks, such as image-based classification [31, 36, 45], object detection [33], and semantic segmentation [41], demonstrating notable efficacy in mitigating catastrophic forgetting. Nevertheless, their focus on unimodal data restricts their applicability in real-world settings, where learning often involves multiple modalities such as vision and language.
Recently, the emergence of Multimodal Large Language Models (MLLMs) such as LLaVA [25, 26] and InternVL [4] has led to tasks and data involving multiple modalities, including vision, language, audio, and so on, thereby stimulating growing interest in Multimodal Continual Learning (MCL) [28, 9, 17, 40]. Compared to traditional continual learning, MCL not only needs to overcome catastrophic forgetting of previously acquired knowledge, but also needs to address cross-modal conflicts and the challenges posed by diverse task formats [9]. These challenges highlight the need for more effective and generalizable approaches to multimodal continual learning, which remain an open and evolving research problem.
Methods UCIT MLLM-DCL
LR Epoch PEFT Parameter setting LR Epoch PEFT Parameter setting
LoRA-FT 2e-4 {1,1,1,1,1,1} LoRA rank = 16 2e-5 {1,3,1,2,1} LoRA rank = 32
O-LoRA 2e-4 {1,1,1,1,1,1} LoRA rank = 96, expert num = 6 2e-5 {1,3,1,2,1} LoRA rank = 160, expert num = 5
MoELoRA 2e-4 {1,1,1,1,1,1} LoRA rank = 18, expert num = 6 2e-5 {1,3,1,2,1} LoRA rank = 35, expert num = 5
ModalPrompt 2e-4 {1,1,1,1,1,1} Prompt prefix len = 10, expert num = 6 2e-4 {1,3,1,2,1} Prompt prefix len = 20, expert num = 5
ModalPrompt* 2e-4 {10,10,10,10,10,10} Prompt prefix len = 10, expert num = 6 2e-4 {10,10,10,10,10} Prompt prefix len = 20, expert num = 5
CL-MoE 2e-4 {1,1,1,1,1,1} LoRA rank = 96, expert num = 6 2e-5 {1,3,1,2,1} LoRA rank = 160, expert num = 5
HiDe 2e-4 {1,1,1,1,1,1} LoRA rank = 96, expert num = 6 2e-5 {1,3,1,2,1} LoRA rank = 160, expert num = 5
SEFE 2e-4 {1,1,1,1,1,1} LoRA rank = 16 2e-5 {1,3,1,2,1} LoRA rank = 32
DISCO 2e-4 {1,1,1,1,1,1} LoRA rank = 96, expert num = 6 2e-5 {1,3,1,2,1} LoRA rank = 160, expert num = 5
Table 1: Training configurations and PEFT settings for all methods. While MoELoRA adopts the parameter extension paradigm, it differs from other methods by not introducing task-specific modules. Its settings are therefore adjusted separately to ensure fair comparison.
In this work, we investigate Multimodal Continual Instruction Tuning (MCIT), a crucial yet less explored task that extends continual learning to the instruction tuning of MLLMs. Under this setting, the model is required to sequentially learn a series of visual question answering tasks that differ significantly in both knowledge domains (e.g. medical and financial) and answer formats (e.g. image captioning and multiple-choice answering), thereby posing greater challenges to existing methods. There are several studies that have introduced benchmarks [2, 44, 7, 22, 43] and methods [42, 27, 15, 3, 39] for MCIT, contributing to the advancement of this field. However, we identify two major limitations in existing works: (1) Overlap between benchmark datasets and the pre-training data of MLLMs. Since MLLM performance heavily depends on large-scale instruction tuning data used in pre-training, many benchmark datasets such as GQA [16] and VQAv2 [6] are often already seen by the models. Using these datasets again can cause information leakage [20] and affect evaluation fairness. (2) Lack of direct comparison between methods under consistent settings. Rapid progress in MCIT has resulted in evaluations conducted on a wide range of benchmarks with varying protocols and settings, which impedes fair and meaningful comparisons among studies and makes it difficult to accurately assess the relative strengths and weaknesses of different methods.
To tackle these challenges and facilitate further research in MCIT, we introduce MCITlib, the first publicly available code library and benchmark for continual instruction tuning of MLLMs. We carefully select benchmarks that avoid information leakage for our experiments and reproduce multiple representative MCIT methods to ensure fair and comprehensive comparisons of their strengths and limitations. MCITlib will be continuously updated to incorporate the latest developments in MCIT, with the goal of supporting and advancing research in the broader field of Multimodal Continual Learning.
2 MCITlib Setup
Implement Algorithms. In MCITlib, we have implemented 8 MCIT algorithms, including LoRA-FT [14], O-LoRA [37], MoELoRA [2], ModalPrompt [42], CL-MoE [15], HiDe-LLaVA [7], SEFE [3], and DISCO [8]. We adopt the commonly used LLaVA-1.5-7b [^1] https://huggingface.co/liuhaotian/llava-v1.5-7b as the base model and employ Parameter-Efficient Fine-Tuning (PEFT) strategies [14, 26] for training. The training process follows the rehearsal-free continual learning setting [46, 47], where data from previous tasks is not reused during the training of new tasks. To ensure a fair comparison, we adopt the original parameter settings of each method as closely as possible, while aligning the PEFT-related configurations. Please refer to Table 1 for detailed settings.
Figure 1: Illustration of the evaluation metric calculations.
Benchmarks and Datasets. Given that the base model has already encountered extensive image-text data during pre-training, we select two benchmarks specifically designed to minimize information leakage during MCIT training:
  • UCIT benchmark [7] consists of six datasets: ArxivQA [23], CLEVR-Math [24], IconQA [30], ImageNet-R [13], VizWiz-Caption [11], and Flickr30k [34]. These datasets include various instruction formats such as image captioning, visual QA, and multiple-choice questions. The base model performs poorly in zero-shot settings, indicating low risk of information leakage. The dataset order is as follows: ImageNet-R $\rightarrow$ ArxivQA $\rightarrow$ VizWiz-Caption $\rightarrow$ IconQA $\rightarrow$ CLEVR-Math $\rightarrow$ Flickr30k.
  • MLLM-DCL benchmark [44] comprises multiple downstream tasks spanning distinct knowledge domains. It includes datasets such as RSVQA [29], PathVQA [12], DriveLM [35], FinVis [38], AI2D [18], Sciverse [10], MapQA [1], and TQA [19], covering five specialized areas: Remote Sensing, Medical, Autonomous Driving, Finance, and Science. The downtask order is as follows: Remote Sensing $\rightarrow$ Medical $\rightarrow$ Autonomous Driving $\rightarrow$ Science $\rightarrow$ Finance.
Method Venue ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k MFT ($\uparrow$) MFN ($\uparrow$) MAA ($\uparrow$) BWT ($\uparrow$)
Zero-shot - 16.27 53.73 38.39 19.20 20.63 41.88 - 31.68 - -
Individual - 91.67 90.83 57.87 78.43 76.63 61.72 - 76.19 - -
LoRA-FT ICLR-22 58.03 77.63 44.39 67.40 61.77 58.22 76.89 61.24 68.57 -18.78
O-LoRA EMNLP-23 77.50 78.07 44.50 63.13 64.73 58.16 76.01 64.35 69.21 -13.99
MoELoRA NeurIPS-24 70.07 77.70 44.69 50.03 54.03 57.34 71.17 58.98 64.74 -14.63
ModalPrompt Arxiv-24 51.07 87.27 48.11 39.23 46.57 42.93 52.65 52.53 68.63 -0.15
ModalPrompt* Arxiv-24 74.43 92.00 55.92 44.27 53.97 43.67 60.76 60.71 62.59 -0.05
CL-MoE CVPR-25 66.33 77.00 44.78 51.87 53.53 57.42 71.46 58.49 63.94 -15.56
HiDE ACL-25 84.03 90.73 44.43 58.93 41.37 54.25 69.96 62.29 65.70 -9.20
SEFE ICML-25 80.83 78.00 47.01 69.63 65.83 57.92 75.98 66.54 70.25 -11.33
DISCO ICCV-25 87.43 93.07 46.96 68.13 65.70 56.69 75.87 69.66 72.71 -7.45
Table 2: Comparison of different methods on the UCIT benchmark. The best performance is shown in bold.
Figure 2: Results of each incremental stage on UCIT benchmark.
Evaluation Metrics. Following the evaluation protocol in SEFE [3], four integrated metrics are used to assess continual learning performance:
  • Mean Finetune Accuracy (MFT) measures the average accuracy achieved on each task immediately after it is learned, serving as an upper bound that reflects the model’s performance in the absence of forgetting.
  • Mean Final Accuracy (MFN) computes the average accuracy over all tasks after completing the full incremental training process, representing the model’s overall retained performance.
  • Mean Average Accuracy (MAA) calculates the mean of average accuracies on all learned tasks after each training step, offering a holistic view of performance throughout the continual learning process.
  • Backward Transfer (BWT) captures the change in accuracy for each task by comparing its final accuracy with that immediately after it was learned, quantifying the extent of forgetting.
For clarity, a conceptual illustration of the four evaluation metrics is provided in Figure 1. Detailed calculation procedures can be found in SEFE [3].
The above features are supported by the current version of MCITlib. In the future, we will continue to track recent advances in the MCL field and expand MCITlib along three key dimensions: benchmarks, methods, and base models. We also welcome community contributions to help build a more comprehensive and impactful resource for Multimodal Continual Learning.
3 Experiment
Method Venue RS Med AD Sci Fin MFT ($\uparrow$) MFN ($\uparrow$) MAA ($\uparrow$) BWT ($\uparrow$)
Zero-shot - 32.29 28.28 15.59 35.55 62.56 - 34.85 - -
Individual - 78.15 58.20 52.77 49.32 88.02 - 65.29 - -
LoRA-FT ICLR-22 69.65 41.59 25.43 40.88 87.45 64.98 53.00 58.52 -14.97
O-LoRA EMNLP-23 74.64 44.42 30.02 41.47 87.15 65.16 55.54 59.53 -12.03
MoELoRA NeurIPS-24 77.54 41.85 27.62 40.13 86.75 64.94 54.78 58.53 -12.71
ModalPrompt Arxiv-24 53.63 45.68 40.77 41.81 87.82 53.87 53.94 53.87 0.09
ModalPrompt* Arxiv-24 78.67 51.37 47.80 43.29 87.78 61.84 61.78 61.81 -0.07
CL-MoE CVPR-25 71.34 46.84 26.33 41.17 88.74 66.06 54.88 59.30 -13.97
HiDE ACL-25 74.31 48.95 33.21 38.54 81.55 60.77 55.31 57.04 -6.82
SEFE ICML-25 77.26 50.37 37.21 40.87 86.82 65.01 58.51 60.96 -8.13
DISCO ICCV-25 76.49 44.48 44.84 46.61 89.22 64.78 60.33 62.41 -5.57
Table 3: Comparison of different methods on the MLLM-DCL benchmark. The best performance is shown in bold.
Figure 3: Results of each incremental stage on MLLM-DCL benchmark.
We report the numerical results of all methods on the two benchmarks in Table 2 and Table 3, while Figure 2 and Figure 3 illustrate their performance curves across training stages. First, we observe that in continual learning tasks for MLLMs, direct sequential fine-tuning does not lead to catastrophic forgetting as severely as in traditional class-incremental learning. This is reflected in the MFN results, where LoRA-FT surpasses the zero-shot baseline by 29.56% and 18.15% on the two benchmarks, respectively. This indicates that the strong inherent generalization ability of MLLMs can partially mitigate forgetting, though a substantial gap remains compared to the upper bound of individual task performance.
Regarding different methods, we observe that there are still notable differences in their ability to alleviate forgetting. Although all methods show improvements in the BWT metric, which suggests some mitigation of forgetting, a closer examination of the MFT shows that certain methods, such as ModalPrompt and HiDe, achieve this by compromising their ability to learn new tasks effectively. In other words, these methods “learn less and forget less”, which ultimately limits overall performance improvement. MFN and MAA reflect the model’s average performance across all tasks and throughout the entire continual learning process, making them more representative. Among all methods, DISCO achieves the best results on both benchmarks. As a parameter extension method, it stores a set of LoRA parameters for each task and selects the appropriate parameter embeddings during inference based on textual similarity, which effectively mitigates forgetting. However, as the number of tasks increases, this approach inevitably leads to considerable parameter overhead. SEFE, as a parameter regularization method, shows strong potential by applying regularizations during new task training to preserve important parameters from previous tasks, thereby mitigating forgetting without adding extra parameters.
The performance curves in Figure 2 and Figure 3 also reveal distinct forgetting patterns across tasks. For example, in the MLLM-DCL benchmark, Task 1 (Remote Sensing) maintained relatively stable performance even after sequentially learning four additional tasks, indicating that various methods were able to mitigate forgetting effectively. In contrast, Task 5 (CLEVR-Math) in the UCIT benchmark experienced significant performance degradation across multiple methods after fine-tuning on only one new task. These observations suggest that MCIT tasks present unique challenges not commonly seen in traditional continual learning scenarios, due to their diverse modalities, task formats, and domain variations, which may lead to more unpredictable or task-specific forgetting patterns.
4 Conclusion
In this paper, we introduce MCITlib, a comprehensive code library designed for continual instruction tuning of Multimodal Large Language Models. The library includes a collection of representative MCIT algorithms and carefully selected benchmarks that reduce information leakage and ensure fair comparisons. By providing unified implementations and evaluation protocols, MCITlib aims to accelerate research progress in Multimodal Continual Learning.
Limitations. The current version of MCITlib is limited in terms of base model diversity and scale, with experiments conducted only on LLaVA-1.5-7B. We have not yet evaluated the continual learning performance of different methods on larger or more diverse MLLMs. In addition, existing metrics mainly focus on benchmark accuracy, while factors such as training/inference efficiency and the impact on the MLLM’s original generalization capabilities are also important. Future updates will address these limitations, extending MCITlib to support more models, tasks, and evaluation dimensions.

### v2

@@
 Individual - 91.67 90.83 57.87 78.43 76.63 61.72 - 76.19 - -
-LoRA-FT ICLR-22 58.03 77.63 44.39 67.40 61.77 58.22 76.89 61.24 68.57 -18.78
-O-LoRA EMNLP-23 77.50 78.07 44.50 63.13 64.73 58.16 76.01 64.35 69.21 -13.99
-MoELoRA NeurIPS-24 70.07 77.70 44.69 50.03 54.03 57.34 71.17 58.98 64.74 -14.63
-ModalPrompt Arxiv-24 51.07 87.27 48.11 39.23 46.57 42.93 52.65 52.53 68.63 -0.15
-ModalPrompt* Arxiv-24 74.43 92.00 55.92 44.27 53.97 43.67 60.76 60.71 62.59 -0.05
-CL-MoE CVPR-25 66.33 77.00 44.78 51.87 53.53 57.42 71.46 58.49 63.94 -15.56
-HiDE ACL-25 84.03 90.73 44.43 58.93 41.37 54.25 69.96 62.29 65.70 -9.20
-SEFE ICML-25 80.83 78.00 47.01 69.63 65.83 57.92 75.98 66.54 70.25 -11.33
-DISCO ICCV-25 87.43 93.07 46.96 68.13 65.70 56.69 75.87 69.66 72.71 -7.45
+LoRA-FT ICLR-22 58.03 77.63 44.39 67.40 61.77 58.22 76.89 61.24 76.55 -18.78
+O-LoRA EMNLP-23 77.50 78.07 44.50 63.13 64.73 58.16 76.01 64.35 78.02 -13.99
+MoELoRA NeurIPS-24 70.07 77.70 44.69 50.03 54.03 57.34 71.17 58.98 75.08 -14.63
+ModalPrompt EMNLP-25 51.07 87.27 48.11 39.23 46.57 42.93 52.65 52.53 57.67 -0.15
+ModalPrompt* EMNLP-25 74.43 92.00 55.92 44.27 53.97 43.67 60.76 60.71 70.57 -0.05
+CL-MoE CVPR-25 66.33 77.00 44.78 51.87 53.53 57.42 71.46 58.49 74.19 -15.56
+HiDE ACL-25 84.03 90.73 44.43 58.93 41.37 54.25 69.96 62.29 77.32 -9.20
+SEFE ICML-25 80.83 78.00 47.01 69.63 65.83 57.92 75.98 66.54 78.76 -11.33
+DISCO ICCV-25 87.43 93.07 46.96 68.13 65.70 56.69 75.87 69.66 81.60 -7.45
 Table 2: Comparison of different methods on the UCIT benchmark. The best performance is shown in bold.
@@
 Individual - 78.15 58.20 52.77 49.32 88.02 - 65.29 - -
-LoRA-FT ICLR-22 69.65 41.59 25.43 40.88 87.45 64.98 53.00 58.52 -14.97
-O-LoRA EMNLP-23 74.64 44.42 30.02 41.47 87.15 65.16 55.54 59.53 -12.03
-MoELoRA NeurIPS-24 77.54 41.85 27.62 40.13 86.75 64.94 54.78 58.53 -12.71
-ModalPrompt Arxiv-24 53.63 45.68 40.77 41.81 87.82 53.87 53.94 53.87 0.09
-ModalPrompt* Arxiv-24 78.67 51.37 47.80 43.29 87.78 61.84 61.78 61.81 -0.07
-CL-MoE CVPR-25 71.34 46.84 26.33 41.17 88.74 66.06 54.88 59.30 -13.97
-HiDE ACL-25 74.31 48.95 33.21 38.54 81.55 60.77 55.31 57.04 -6.82
-SEFE ICML-25 77.26 50.37 37.21 40.87 86.82 65.01 58.51 60.96 -8.13
-DISCO ICCV-25 76.49 44.48 44.84 46.61 89.22 64.78 60.33 62.41 -5.57
+LoRA-FT ICLR-22 69.65 41.59 25.43 40.88 87.45 64.98 53.00 61.13 -14.97
+O-LoRA EMNLP-23 74.64 44.42 30.02 41.47 87.15 65.16 55.54 62.12 -12.03
+MoELoRA NeurIPS-24 77.54 41.85 27.62 40.13 86.75 64.94 54.78 61.76 -12.71
+ModalPrompt EMNLP-25 53.63 45.68 40.77 41.81 87.82 53.87 53.94 49.67 0.09
+ModalPrompt* EMNLP-25 78.67 51.37 47.80 43.29 87.78 61.84 61.78 64.05 -0.07
+CL-MoE CVPR-25 71.34 46.84 26.33 41.17 88.74 66.06 54.88 61.79 -13.97
+HiDE ACL-25 74.31 48.95 33.21 38.54 81.55 60.77 55.31 60.68 -6.82
+SEFE ICML-25 77.26 50.37 37.21 40.87 86.82 65.01 58.51 63.63 -8.13
+DISCO ICCV-25 76.49 44.48 44.84 46.61 89.22 64.78 60.33 63.93 -5.57
 Table 3: Comparison of different methods on the MLLM-DCL benchmark. The best performance is shown in bold.
 Figure 3: Results of each incremental stage on MLLM-DCL benchmark.
-We report the numerical results of all methods on the two benchmarks in Table 2 and Table 3, while Figure 2 and Figure 3 illustrate their performance curves across training stages.
+We present the numerical results of all methods on the two benchmarks in Table 2 and Table 3.
+Meanwhile, Figure 2 and Figure 3 illustrate the performance curves across training stages.
+The complete results matrix for each method is provided in Appendix A.
 First, we observe that in continual learning tasks for MLLMs, direct sequential fine-tuning does not lead to catastrophic forgetting as severely as in traditional class-incremental learning.
@@
 Future updates will address these limitations, extending MCITlib to support more models, tasks, and evaluation dimensions.
+Supplementary Material
+Appendix A Detailed Continual Learning Results
+In this section, we show the detailed results for all methods at each stage of continual learning on LLaVA-1.5-7b.
+ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
+ImgNet-R 92.00
+ArxivQA 90.63 91.03
+VizWiz 74.00 91.4 61.9
+IconQA 72.73 78.53 53.03 79.87
+CLEVR 68.97 77.6 49.43 67.67 78.33
+Flickr30k 58.03 77.63 44.39 67.4 61.77 58.22
+Table 4: Result matrix of LoRA-FT on UCIT benchmark.
+ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
+ImgNet-R 91.40
+ArxivQA 90.07 94.73
+VizWiz 81.50 91.70 61.59
+IconQA 82.4 77.73 54.06 76.63
+CLEVR 81.07 77.87 50.62 62.00 73.53
+Flickr30k 77.50 78.07 44.50 63.13 64.73 58.16
+Table 5: Result matrix of O-LoRA on UCIT benchmark.
+ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
+ImgNet-R 91.23
+ArxivQA 90.07 92.80
+VizWiz 80.30 92.03 60.22
+IconQA 80.53 77.20 51.19 53.73
+CLEVR 79.37 77.40 48.68 51.30 71.67
+Flickr30k 70.07 77.70 44.69 50.03 54.03 57.34
+Table 6: Result matrix of MoELoRA on UCIT benchmark.
+ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
+ImgNet-R 51.10
+ArxivQA 51.03 87.37
+VizWiz 51.10 87.37 48.14
+IconQA 51.20 87.27 48.03 39.30
+CLEVR 51.17 87.23 47.96 39.40 47.07
+Flickr30k 51.07 87.27 48.11 39.23 46.57 42.93
+Table 7: Result matrix of ModalPrompt on UCIT benchmark.
+ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
+ImgNet-R 74.57
+ArxivQA 74.73 91.83
+VizWiz 74.53 91.93 55.76
+IconQA 74.53 91.90 55.97 44.20
+CLEVR 74.50 91.83 55.75 44.23 54.50
+Flickr30k 74.43 92.00 55.92 44.27 53.97 43.67
+Table 8: Result matrix of ModalPrompt* on UCIT benchmark.
+ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
+ImgNet-R 91.30
+ArxivQA 87.40 92.27
+VizWiz 76.23 90.53 60.18
+IconQA 77.33 76.20 51.09 54.33
+CLEVR 75.73 75.70 48.27 52.80 73.23
+Flickr30k 66.33 77.00 44.78 51.87 53.53 57.42
+Table 9: Result matrix of CL-MoE on UCIT benchmark.
+ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
+ImgNet-R 90.87
+ArxivQA 88.83 92.27
+VizWiz 88.17 91.90 54.46
+IconQA 87.00 89.70 46.51 73.13
+CLEVR 86.23 90.07 44.28 64.37 54.77
+Flickr30k 84.03 90.73 44.43 58.93 41.37 54.25
+Table 10: Result matrix of HiDe on UCIT benchmark.
+ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
+ImgNet-R 90.80
+ArxivQA 90.23 91.23
+VizWiz 83.60 89.73 61.24
+IconQA 84.17 77.20 56.36 78.93
+CLEVR 82.90 77.13 53.54 71.43 75.73
+Flickr30k 80.83 78.00 47.01 69.63 65.83 57.92
+Table 11: Result matrix of SEFE on UCIT benchmark.
+ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
+ImgNet-R 90.97
+ArxivQA 90.07 91.77
+VizWiz 89.40 92.53 61.39
+IconQA 88.50 93.97 58.61 78.00
+CLEVR 88.03 93.40 57.11 70.93 76.40
+Flickr30k 87.43 93.07 46.96 68.13 65.70 56.69
+Table 12: Result matrix of DISCO on UCIT benchmark.
+RS Med AD Sci Fin
+RS 78.32
+Med 74.68 57.53
+AD 68.93 47.19 52.15
+Sci 75.12 45.56 38.46 49.44
+Fin 69.65 41.59 25.43 40.88 87.45
+Table 13: Result matrix of LoRA-FT on DCL benchmark.
+RS Med AD Sci Fin
+RS 79.25
+Med 74.05 56.52
+AD 76.06 43.71 52.32
+Sci 76.60 44.87 40.57 50.58
+Fin 74.64 44.42 30.02 41.47 87.15
+Table 14: Result matrix of O-LoRA on DCL benchmark.
+RS Med AD Sci Fin
+RS 79.09
+Med 74.78 58.73
+AD 77.69 43.72 51.47
+Sci 76.87 43.79 32.81 48.67
+Fin 77.54 41.85 27.62 40.13 86.75
+Table 15: Result matrix of MoELoRA on DCL benchmark.
+RS Med AD Sci Fin
+RS 53.12
+Med 53.02 45.82
+AD 53.14 45.77 40.78
+Sci 53.03 45.65 40.78 41.83
+Fin 53.63 45.68 40.77 41.81 87.82
+Table 16: Result matrix of ModalPrompt on DCL benchmark.
+RS Med AD Sci Fin
+RS 78.76
+Med 78.57 51.46
+AD 78.61 51.59 47.92
+Sci 78.59 51.50 47.90 43.28
+Fin 78.67 51.37 47.80 43.29 87.78
+Table 17: Result matrix of ModalPrompt* on DCL benchmark.
+RS Med AD Sci Fin
+RS 78.76
+Med 78.57 51.46
+AD 78.61 51.59 47.92
+Sci 78.59 51.50 47.90 43.28
+Fin 78.67 51.37 47.80 43.29 87.78
+Table 18: Result matrix of ModalPrompt* on DCL benchmark.
+RS Med AD Sci Fin
+RS 79.08
+Med 73.48 60.56
+AD 72.61 44.42 51.62
+Sci 71.02 48.04 37.70 50.28
+Fin 71.34 46.84 26.33 41.17 88.74
+Table 19: Result matrix of CL-MoE on DCL benchmark.
+RS Med AD Sci Fin
+RS 78.14
+Med 74.26 58.05
+AD 74.90 42.94 39.65
+Sci 75.43 44.91 38.33 46.44
+Fin 74.31 48.95 33.21 38.54 81.55
+Table 20: Result matrix of HiDe on DCL benchmark.
+RS Med AD Sci Fin
+RS 78.27
+Med 76.32 58.42
+AD 77.22 49.13 52.49
+Sci 77.83 47.7 43.01 49.04
+Fin 77.26 50.37 37.21 40.87 86.82
+Table 21: Result matrix of SEFE on DCL benchmark.
+RS Med AD Sci Fin
+RS 78.57
+Med 75.80 52.36
+AD 76.37 49.78 53.04
+Sci 76.11 45.76 49.26 49.23
+Fin 76.03 45.20 43.79 42.33 88.95
+Table 22: Result matrix of DISCO on DCL benchmark.